OpenStudy (anonymous):

Solve for n. A. n = 6 B. n = 5 C. n = 30 D. n = 75

OpenStudy (mathmale):

Solve what for n? ;)

OpenStudy (anonymous):

n/18 = 5/15

OpenStudy (anonymous):

@mathmale @LocalHero7

OpenStudy (anonymous):

@mathmale dont leaveeee meeseeeeee xD

OpenStudy (anonymous):

use the cross multiplication method

OpenStudy (anonymous):

the result would be n(15) = 18(5)

OpenStudy (anonymous):

90?

OpenStudy (kayders1997):

Yes do that

OpenStudy (anonymous):

no, you forgot to divide both sides by 15

OpenStudy (kayders1997):

90= what though

OpenStudy (anonymous):

ohhh 6??!

OpenStudy (kayders1997):

Yes!

OpenStudy (anonymous):

yay im sooooo smart

OpenStudy (anonymous):

jk lol thanks guysssss

OpenStudy (anonymous):

np :)

OpenStudy (kayders1997):

You are smart!

OpenStudy (kayders1997):

Good job :)